

********************************************************************
UARS WEEKLY STATUS REPORT - 09 AUG 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off for the yaw-around on 7/31 and was repowered on
8/5. Data taking resumed about 40 hours after power on because the
instrument was too cold. Otherwise, SUSIM successfully performed its
daily mid-resolution and MgII solar scans during periods of predicted
UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from this conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes. A PEM representative is attending
the TIMED workshop August 7-9.
PEM data from the second GUVI-PEM conjunction is now being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII experienced difficulties with the new ECLIPSE commanding
system, which appeared to be resolved, in that WINDII operations
were successfully achieved. However, we subsequently learned
that no science data were being produced. For the next turn-on,
scheduled August 10, we have planned tests to help determine
whether this problem originates with the uploads normally sent.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission. The UARS Monthly Status Review
Meeting was held on August 07, 2002.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 July 2002 (Orbit 59468, GMT Day 208) through 02
August 2002 (Orbit 59571, GMT Day 214).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 210/1643 59507 POWER ON / DAY ONLY
HRDI 210/1824 59508 DAY ONLY TO LOW PWR
HALOE 210/2157 59511 POWER OFF
HRDI 210/2341 59512 POWER ON / DAY ONLY
PEM 211/2156 59526 POWER OFF
SOLSTICE 212/2155 59541 POWER OFF
SUSIM 212/2155 59541 POWER OFF
A Forward-to-Reverse Yaw Around maneuver (number 111) was successfully
executed in Orbit 59566 beginning 214/14:13:17Z and completing at
214/14:50:27Z on 02 August 2002. All instruments were in their off or
standby mode (including HRDI and PEM). Due to the SA position at 269
degrees, the effective night length of the spacecraft after the
maneuver was approximately 57.07 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus
voltage was 25.14 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 20.51 and 20.16 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 04
September 2002 (DOY 247).
The beta angle ranged between the angles of 19.4 degrees to 0.0
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 45 degrees on 19 August 2002 (DOY 231).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.59 to 1.39 degrees C. Additional
battery performance data for each battery on 26 July and 02 August
2002 are provided in Appendix D.
The clock error ranged from 9.9 to 10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HRDI = 03
HALOE = 01
PEM = 01
SOLSTICE = 01
SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Raytheon development team has delivered a new database patch which
should fix the status buffer display problem. This patch will be
tested this week on the prototype system.
The FOT is also working with the WINDII instrument team to determine
the source of their instrument turn on problem. It was recently
discovered, by the WINDII team, that WINDII's science data has not
been captured correctly since February. There were discrepancies
found in their POCC resident microloads during Eclipse transition and
the investigation has been focused on determining if these loads are
correct.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 04 September 2002 (DOY 247).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT
Occurrences Appendix C, Clock & Solar Array Rate
Corrections Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 August 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
59507 210/1643 02-259 HRDI POWER ON / DAY ONLY
59508 210/1824 02-260 HRDI DAY ONLY TO LOW POWER
59511 210/2157 02-261 HALOE POWER OFF
59512 210/2341 02-262 HRDI POWER ON / DAY ONLY
59526 211/2156 02-263 PEM POWER OFF
59527 211/2340 02-264 OBC/CLK CLK ADJ FR 0.00 TO 20.6 MSEC/DAY
59541 212/2155 02-265 SOLSTICE POWER OFF
59541 212/2155 02-266 SUSIM POWER OFF
59546 213/0626 02-268 OBC/CLK CLK ADJ FR 20.6 TO 0.00 MSEC/DAY
59565 214/1234 02-268 OBC YAW FWD TO REV
59566 214/1439 02-268 OBC YAW
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 August 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 August 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
59527 211/2345 CLK ADJ FR 0.00 TO 20.6 MSEC/DAY
59546 213/0626 CLK ADJ FR 20.6 TO 0.00 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 August 2002
26 July 2002 - GMT Day 207 Beta = 28.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -157.0/-168.0 -3.30/-4.24 +00.0/+00.0 08.5 0.000 0.0
2 + 11.2/- 69.6 +6.44/+5.15 +27.5/-17.9 27.8 1.030 21.8
3 + 22.4/- 56.0 +5.21/+3.87 +26.7/-17.6 27.7 1.020 21.4
02 August 2002 - GMT Day 214 Beta = 0.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -157.0/-168.0 -3.05/-4.62 +00.0/+00.0 08.5 0.000 0.0
2 + 11.2/- 69.6 +5.47/+4.08 +27.0/-13.4 23.8 1.090 20.4
3 + 22.4/- 50.4 +4.51/+2.92 +25.7/-13.5 23.8 1.080 20.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
59565 214/1238 L5 -> L6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 August 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-042 FOT 59503 210/0959 N/A TMON 01 tripped
02-039 FOT 59526 211/2156 30832 Generic Late Acquisition
02-043 FOT 59571 214/2308 31245 Data Drop Out
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT See List July 30832 Generic Late Acquisition
02-041 FOT 59279 195/1156 N/A TMON 01 tripped
02-042 FOT 59503 210/0959 N/A TMON 01 tripped
02-043 FOT 59571 214/2308 31245 Data Drop Out